I was around the 1Y1 rack today. Trials were done to get the PDFR of AS55.
We woke up the PDFR measurement setup that has been sleeping since summer. We ran a check for the laser module and the multiplexer module. We tried setting things up for measuring frequency response of AS55.
We could not repeat Nichin's measurements because the gpib scripts are outdated and need to be revised.
PDFR diode laser was shutdown after this job.